# Repetitive Contrastive Learning Enhances Mamba’s Selectivity in Time Series Prediction

### v1

@@ 4.7 Comparison of Replacement and Freezing Methods
up-rate-1.0489 -4.7865 0.7322 -4.1605 4.8065 4.5439 8.1034 12.4834
Table 8: Comparison of Replacement and Freezing Methods. The ”layer-x%” indicates that the first xof layers were replaced by pre-trained blocks.
@@ 5.2 Analysis of Enhanced Selectivity
We compared the Focus Ratio and Memory Entropy of the Mamba block when modeling time series with and without RCL, and the results are shown in Table 6. It can be observed that after applying RCL, the Focus Ratio significantly improves, indicating more pronounced processes of significant memory and significant forgetting. This suggests that the model becomes more focused on key information and more decisive in ignoring noisy information. Similarly, RCL also leads to a notable increase in Memory Entropy, demonstrating that the Mamba block’s memory patterns for time-step information become more complex and diverse. This enables the model to better capture essential aspects of the sequence with greater selectivity.
@@
5.3 Theoretical Analysis of Repetitive Contrastive Learning
@@
Minimizing the InfoNCE loss can be interpreted as maximizing the lower bound of mutual information between the anchor and the positive example(van den Oord et al., 2019; Wu et al., 2020):
@@
$-\log\left(1+\exp\left(\frac{x_{t+1}^{2}-x_{t+1}(x_{t}+\sigma)}{(x_{t}+\sigma)^{2}}h_{t}^{2}\right)\right)$
@@
$-\log\left(1+\exp\left(\frac{h_{t}^{2}x_{t+1}^{4}}{(x_{t}+\sigma)^{3}}+\frac{h_{t}^{2}}{x_{t}+\sigma}\right)\right)$
@@
6 Visualization of Comparative Effects
6.1 Visualization of Embedding in Augmentation Sequence
To visually demonstrate the impact of our contrastive learning methods, we plotted the cosine similarity values(Rahutomo et al., 2012) between embedding vectors of the same input sequence from the ETTm1 dataset using a heatmap (Shin et al., 2006). This comparison involves identical Mamba blocks—one trained without contrastive pre-training and the other with it. The resulting variations in distribution highlight the influence of our pre-training objectives, which enhance the model’s ability to selectively focus on relevant features. The images illustrate the differences in the embedding space (Figure 7) and the refined distribution achieved through contrastive learning (Figure 7).
Figure 6: Embedding with contrastive pre-training result
Figure 7: Embedding without pre-training result
6.2 Visualization of clustering of Positive and Negative Cases
We also visualized the detailed distribution of vectors using the UMAP technique for dimensionality reduction, where the original dimensionality of the embedding vectors is 32. UMAP is based on a theoretical framework rooted in Riemannian geometry and algebraic topology, resulting in a scalable and practical algorithm suitable for contrastive learning data (McInnes et al., 2020). In the visualizations (Figure 8), we randomly selected embedding vectors from input sequences and plotted the corresponding vectors for both positive and negative pairs in our method.
7 Related Work
7.1 Models in Deep Time Series Forecasting
Extensive research has been conducted to address time series forecasting problems, primarily focused on proposing new models that improve prediction accuracy. These models can be categorized into five primary groups: Transformer-based, RNN-basedHochreiter and Schmidhuber (1997), CNN-based, MLP-based, and Mamba-based. While emphasizing different aspects, these approaches aim to address key challenges of time series tasks.
Some MLP-based models, such as N-BeatsOreshkin et al. (2020) and N-HiTsChallu et al. (2022), utilize basis approximation and residual connections. TimesNet Wu et al. (2023), a CNN-based model, employs periodical segmentations in the frequency and time domains, extracting inter-period and intra-period patterns. TimeMixer Wang et al. (2024), built solely with MLP and pooling layers, excels by decomposing and mixing multi-scale data.
Transformer-based models like LogTransLi et al. (2020), InformerZhou et al. (2021), AutoformerWu et al. (2022), and FEDformerZhou et al. (2022) enhance adaptability using sparse attention and decomposition techniques. PatchTSTNie et al. (2023) segments time series into patches for denoising, while iTransformerLiu et al. (2024) redefines time embeddings. Mamba-based models, like TimeMachineAhamed and Cheng (2024), unify channel-mixing and independence to refine content selection.
7.2 Contrastive Learning
Most contrastive self-supervised learning methods have been applied in vision Jaiswal et al. (2021) and multimodal learning Manzoor et al. (2024), leveraging high-level attributes that are easily distinguishable and less affected by noise. For example, images remain interpretable despite perturbations like color changes or geometric transformations, while multimodal methods enhance contrast by using cross-modality correlations, such as visual-textual pairing.
In contrast, applying contrastive learning to unimodal sequential data is less common and often requires tailored features. For instance, CodeRetriever Li et al. (2022) employs similarity contrastive loss to capture nuances in code sequences. Sequential recommendation Xie and Li (2024) and text summarization Xu et al. (2022) rely on specialized sequence representations and training techniques.
In time series, contrastive pre-training has improved representation learning. TS2Vec Yue et al. (2022) introduced a universal framework using context view augmentation and hierarchical contrastive learning. TF-C Zhang et al. (2022b) aligned time-based and frequency-based representations for better performance. InfoTS applied information theory to prioritize diverse and high-fidelity representations, while SoftCLT Lee et al. (2024) captured inter-sample and intra-temporal relationships through soft assignments.
8 Conclusion
In this paper, we introduce Repetitive Contrastive Learning (RCL), a novel training paradigm designed to enhance the selective capabilities of Mamba blocks and enable the transfer of these parameters to various Mamba-based backbone models, improving their performance. RCL combines sequence repetition with intra-sequence and inter-sequence contrastive learning, strengthening Mamba blocks’ ability to retain critical information and filter out noise. Through extensive experiments, we demonstrate RCL’s effectiveness across multiple Mamba-based backbone models, significantly boosting their temporal prediction capabilities. From theoretical, qualitative, and quantitative perspectives, we validate the enhanced selective performance achieved by RCL and confirm that it adds no extra memory overhead. In future work, RCL could be adapted to other models with similar principles, such as RNNs and vanilla attention mechanisms.
Acknowledgments
Appendix A Reproducibility Statement
We provide simplified code available at this Anonymous Github link [^1] https://anonymous.4open.science/r/PretrainMamba-DD5B/. You can use this code to reproduce our results by referring to the parameters outlined in the paper.
Appendix B Additional Experimental Results
B.1 Comprehensive Results of Comparison with Temporal Model
@@
B.2 Detail Comparison of Improvements
To demonstrate that pre-training Mamba blocks with RCL can effectively enhance the temporal prediction capabilities of Mamba-based models, we present the performance improvements of four Mamba-based models after using pre-trained parameters. We conducted extensive testing on six datasets, each with an input length of 96 and prediction lengths of $\{96,192,336,720\}$. To clearly illustrate the performance improvements, we provide the percentage increase in MSE and MAE when using pre-trained parameters compared to not using them, as shown by the up-rate in Table 12.

### v2

@@ 4.7 Comparison of Replacement and Freezing Methods
up-rate% -1.0489 -4.7865 0.7322 -4.1605 4.8065 4.5439 8.1034 12.4834
Table 8: Comparison of Replacement and Freezing Methods. The ”layer-x%” indicates that the first x% of layers were replaced by pre-trained blocks.
@@
Figure 4: Visualization of Predictions when Replacing 100%/50% of the Layers and Frozen/Not Forzen Parameters
We visualized the predictions under different settings of replacing 100%/50% of the parameters and whether to freeze them, as shown in Fig. 4, providing a more intuitive understanding of the impact of these configurations. Overall, the performance of replacing 100% + Not Frozen is similar to replacing 50% + Frozen, while replacing 100% + Frozen and replacing 50% + Not Frozen show poorer performance. This is because, for multi-layer models (whether Mamba or Transformer), deeper layers tend to model specificity, while shallower layers tend to model commonality. For time series, this can be intuitively understood as short-term fluctuations and long-term periodicity. From the practical results, comparing (a) and (b), replacing only 50% retains sufficient parameters to focus on learning specific predictive features, whereas replacing 100% leads to poorer specificity fitting and reduced ability to fit extreme values. Comparing (b) and (c), not freezing any parameters alleviates the issue of insufficient trainable parameters when replacing 100%, achieving comparable performance. Comparing (a) and (d), when replacing fewer layers and not freezing parameters, the information learned by RCL can be lost, thus requiring freezing the matrix to retain the unique information learned during pre-training.
4.8 Few-Shot Forecasting
Setup To verify the performance of our method in few-shot scenarios, we designed the Few-Shot Forecasting experiment. Specifically, we retained only the first 10% of the original dataset as the training set to pre-train the Mamba block and train the backbone model, while keeping the size of the test set unchanged. In this setup, not only is the amount of training data significantly reduced, but there is also a substantial temporal gap between the test set and the training set, further increasing the difficulty of the prediction task.
Model TimeMachine* TimeMachine Bi-Mamba* Bi-Mamba iTransformer TimeMixer CrossFormer TimesNet Informer
Metric MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE
ETTh1 0.452 0.471 0.479 0.504 0.504 0.476 0.464 0.489 0.470 0.499 0.728 1.025 0.800 1.405 0.863 1.345 0.814 1.164
ETTh2 0.353 0.319 0.360 0.325 0.325 0.320 0.369 0.335 0.363 0.323 1.165 2.692 1.216 2.588 1.976 5.966 1.149 2.264
ETTm1 0.403 0.419 0.423 0.438 0.438 0.426 0.417 0.438 0.428 0.442 0.551 0.616 0.631 0.846 1.057 1.960 0.821 1.125
ETTm2 0.241 0.171 0.263 0.188 0.239 0.185 0.266 0.191 0.271 0.194 0.400 0.315 0.896 1.497 1.347 3.058 1.284 2.763
Traffic 0.331 0.517 0.341 0.534 0.328 0.493 0.335 0.505 0.332 0.500 0.566 1.040 0.463 0.892 0.443 0.888 0.754 1.453
Electricity 0.259 0.193 0.277 0.199 0.251 0.181 0.268 0.189 0.264 0.188 0.409 0.333 0.363 0.278 0.600 0.707 0.871 1.255
Table 9: Comparative Results Under Few-Shot Setting. The best results are bolded, and the second-best results are underlined.
Result The comparative results are shown in Table 9. TimeMachine* and Bi-Mamba* represent the RCL-enhanced versions of TimeMachine and Bi-Mamba, respectively. First, in the few-shot forecasting problem, the performance of all backbone models experienced a significant decline, indicating that the backbone models are heavily dependent on the amount of data. However, by comparing the results of TimeMachine and Bi-Mamba with and without RCL, it is evident that our method significantly enhances the performance of the backbone models, bringing their capabilities closer to the results achieved with the full dataset. This demonstrates that RCL can uncover more potential temporal patterns even with limited data, thereby reducing the backbone models’ reliance on the volume of data.
4.9 Long-Term Forecasting Under Missing Points Data
Setup To validate the performance of our method under the condition of missing points, we conducted experiments with missing data. Specifically, we randomly masked 10% of the input sequence and predicted future variations under the scenario of missing data. Both the RCL and backbone model training utilized the same masking results. This experimental setup allows us to verify the performance of our method in the presence of missing data.
Model TimeMachine* TimeMachine Bi-Mamba* Bi-Mamba iTransformer TimeMixer CrossFormer TimesNet Informer
Metric MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE
ETTh1 0.391 0.389 0.397 0.397 0.399 0.401 0.404 0.412 0.393 0.386 0.509 0.621 0.408 0.418 0.648 0.716 0.636 0.856
ETTh2 0.321 0.279 0.333 0.288 0.329 0.286 0.335 0.291 0.343 0.305 0.397 0.360 0.516 0.531 0.814 1.129 1.171 2.200
ETTm1 0.352 0.324 0.361 0.344 0.349 0.318 0.354 0.329 0.361 0.344 0.368 0.354 0.430 0.397 0.538 0.537 0.608 0.690
ETTm2 0.253 0.176 0.260 0.182 0.256 0.177 0.261 0.183 0.257 0.177 0.278 0.196 0.347 0.260 0.548 0.526 0.581 0.551
Traffic 0.308 0.451 0.310 0.456 0.308 0.460 0.310 0.462 0.321 0.468 0.420 0.699 0.309 0.583 0.337 0.693 0.438 0.832
Electricity 0.262 0.171 0.268 0.177 0.265 0.179 0.273 0.186 0.272 0.189 0.321 0.236 0.309 0.215 0.473 0.461 0.541 0.546
Table 10: Comparative Results Under Missing Points Data. The best results are bolded, and the second-best results are underlined.
Result The comparative results are shown in Table 10. TimeMachine* and Bi-Mamba* represent the RCL-enhanced versions of TimeMachine and Bi-Mamba, respectively. In the case of missing data points, the backbone models TimeMachine and Bi-Mamba performed worse than other backbone models, such as iTransformer. With the enhancement of RCL, the performance of TimeMachine* and Bi-Mamba* significantly improved, surpassing other models. This demonstrates that RCL can still bring significant improvements to Mamba-based models even under the condition of missing points.
@@ 5.2 Analysis of Enhanced Selectivity
We compared the Focus Ratio and Memory Entropy of the Mamba block when modeling time series with and without RCL, and the results are shown in Table 7. The statistical significance test and correlation analysis of Focus Ratio (FR) and Memory Entropy (ME) are presented in D.2. It can be observed that after applying RCL, the Focus Ratio significantly improves, indicating more pronounced processes of significant memory and significant forgetting. This suggests that the model becomes more focused on key information and more decisive in ignoring noisy information. Similarly, RCL also leads to a notable increase in Memory Entropy, demonstrating that the Mamba block’s memory patterns for time-step information become more complex and diverse. This enables the model to better capture essential aspects of the sequence with greater selectivity.
@@
6 Visualization of Comparative Effects
6.1 Visualization of Embedding in Augmentation Sequence
To visually demonstrate the impact of our contrastive learning methods, we plotted the cosine similarity values(inproceedings) between embedding vectors of the same input sequence from the ETTm1 dataset using a heatmap (JSSv016c03). This comparison involves identical Mamba blocks—one trained without contrastive pre-training and the other with it. The resulting variations in distribution highlight the influence of our pre-training objectives, which enhance the model’s ability to selectively focus on relevant features. The images illustrate the differences in the embedding space (Figure 8) and the refined distribution achieved through contrastive learning (Figure 8).
Figure 7: Embedding with contrastive pre-training result
Figure 8: Embedding without pre-training result
6.2 Visualization of clustering of Positive and Negative Cases
We also visualized the detailed distribution of vectors using the UMAP technique for dimensionality reduction, where the original dimensionality of the embedding vectors is 32. UMAP is based on a theoretical framework rooted in Riemannian geometry and algebraic topology, resulting in a scalable and practical algorithm suitable for contrastive learning data (mcinnes2020umapuniformmanifoldapproximation). In the visualizations (Figure 9), we randomly selected embedding vectors from input sequences and plotted the corresponding vectors for both positive and negative pairs in our method.
6.3 Case Study
Figure 10: Visualization of prediction results with and without RCL across various datasets.
We conducted a case study on multiple datasets for both with RCL and without RCL scenarios, and the visualization results are shown in Fig, 10. The backbone model selected is Timemachine. It can be observed that after incorporating RCL, the model is able to make more accurate predictions, with the predicted curve (red) closely aligning with the actual curve (yellow). More importantly, with RCL, the backbone model can make more accurate predictions and adapt more quickly to extreme cases. As highlighted in Fig, 10, the predicted values with RCL are significantly closer to the true values in extreme scenarios. This demonstrates that RCL is a highly effective method, capable of adapting to extreme changes in time series patterns through enhanced memory and forgetting mechanisms for extreme cases.
7 Related Work
7.1 Models in Deep Time Series Forecasting
Extensive research has been conducted to address time series forecasting problems, primarily focused on proposing new models that improve prediction accuracy. These models can be categorized into five primary groups: Transformer-based, RNN-based10.1162/neco.1997.9.8.1735, CNN-based, MLP-based, Mamba-based and spatio-temporal models which use extra graph data. While emphasizing different aspects, these approaches aim to address key challenges of time series tasks.
Some MLP-based models, such as N-Beatsoreshkin2020nbeatsneuralbasisexpansion and N-HiTschallu2022nhitsneuralhierarchicalinterpolation, utilize basis approximation and residual connections. TimesNet wu2023timesnettemporal2dvariationmodeling, a CNN-based model, employs periodical segmentations in the frequency and time domains, extracting inter-period and intra-period patterns. TimeMixer wang2024timemixerdecomposablemultiscalemixing, built solely with MLP and pooling layers, excels by decomposing and mixing multi-scale data.
Transformer-based models like LogTransli2020enhancinglocalitybreakingmemory, Informerzhou2021informerefficienttransformerlong, Autoformerwu2022autoformerdecompositiontransformersautocorrelation, TCGPNyan2024tcgpn, DPA-STIFormeryan2024double and FEDformerzhou2022fedformerfrequencyenhanceddecomposed enhance adaptability using sparse attention and decomposition techniques. PatchTSTnie2023timeseriesworth64 segments time series into patches for denoising, while iTransformerliu2024itransformerinvertedtransformerseffective redefines time embeddings.
Mamba-based models like TimeMachineahamed2024timemachinetimeseriesworth, HIGSTMyan2025hierarchical combine channel-mixing and independence to enhance content selection, while S-Mambawang2024mambaeffectivetimeseries achieves transformer-like generalization with reduced computational resources.
Spatio-temporal models like STGNN ALIGHTWEIGHTdecompose time series into seasonal and trend components using downsampling and patched sampling, while SAMBAMAMBAFINANCIAL employs bidirectional Mamba blocks and adaptive graph convolution to model long-term dependencies and spatial stock feature interactions with near-linear complexity, and DSTGCN10770607 integrates graph convolution with spatial attention to efficiently capture spatial relationships.
7.2 Contrastive Learning
Most contrastive self-supervised learning methods have been applied in vision jaiswal2021surveycontrastiveselfsupervisedlearning and multimodal learning manzoor2024multimodalityrepresentationlearningsurvey, leveraging high-level attributes that are easily distinguishable and less affected by noise. For example, images remain interpretable despite perturbations like color changes or geometric transformations, while multimodal methods enhance contrast by using cross-modality correlations, such as visual-textual pairing.
In contrast, applying contrastive learning to unimodal sequential data is less common and often requires tailored features. For instance, CodeRetriever li2022coderetrieverunimodalbimodalcontrastive employs similarity contrastive loss to capture nuances in code sequences. ECP-MambaECPMAMBA integrate multi-scale self-supervised contrastive learning with a state space model (SSM) backbone in image modality. Sequential recommendation XIE2024112257 and text summarization xu2022sequencelevelcontrastivelearning rely on specialized sequence representations and training techniques.
In time series, contrastive pre-training has improved representation learning. TS2Vec yue2022ts2vecuniversalrepresentationtime introduced a universal framework using context view augmentation and hierarchical contrastive learning. TF-C zhang2022selfsupervisedcontrastivepretrainingtime aligned time-based and frequency-based representations for better performance. InfoTS applied information theory to prioritize diverse and high-fidelity representations, while SoftCLT softcls captured inter-sample and intra-temporal relationships through soft assignments.
8 Conclusion
In this paper, we introduce Repetitive Contrastive Learning (RCL), a novel training paradigm designed to enhance the selective capabilities of Mamba blocks and enable the transfer of these parameters to various Mamba-based backbone models, improving their performance. RCL combines sequence repetition with intra-sequence and inter-sequence contrastive learning, strengthening Mamba blocks’ ability to retain critical information and filter out noise. Through extensive experiments, we demonstrate RCL’s effectiveness across multiple Mamba-based backbone models and diverse temporal prediction tasks, significantly enhancing their temporal modeling and forecasting capabilities. From theoretical, qualitative, and quantitative perspectives, we validate the enhanced selective performance achieved by RCL and confirm that it adds no extra memory overhead. In addition, we have validated the generalization capability of RCL across other types of models. Through experiments, we have demonstrated that RCL is significantly effective on step-by-step models like RNNs, and we have analyzed why it is not well-suited for non-step-by-step models such as Transformers.
Acknowledgments
Appendix A Reproducibility Statement
We provide simplified code available at this Github link [^1] https://github.com/xiaxiaoguang/PretrainMamba. You can use this code to reproduce our results by referring to the parameters outlined in the paper.
Appendix B Algorithm Workflow
[…]
Algorithm 1 Repetitive Contrastive Learning
1:Input $\mathbf{X}$: $(B,T,F)$
2:Outputs $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$
[…]
4:for $j=1$ to $n_{t}-1$ do
5:  Sample $Noise_{j}\sim\mathcal{N}(0,\sigma_{j}^{2})$: $(B,T,F)$
6:end for
[…]
8:NoiseNoise: $(B,T,n_{t},F)$ $\leftarrow$ Stack $Noise_{0},\dots,Noise_{n_{t}-1}]$
[…]
14:return $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$
In the second stage, an arbitrary Mamba-based backbone model is selected, and the initialization parameters of its Mamba blocks are replaced with $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$. The algorithm is illustrated in Algorithm 2.
Algorithm 2 Replace and Train Backbone Model
Input $\mathbf{X}$: $(B,T,F)$, model $\mathbb{M}$, parameters $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$
2:Trained model $\mathbb{M}^{\prime}_{aug}$
$\mathbb{M}_{aug}$: Replace Mamba Block‘s parameters in $\mathbb{M}$ with $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$
[…]
return $M^{\prime}_{aug}$
Appendix C Theoretical Analysis of Repetitive Contrastive Learning
@@
Minimizing the InfoNCE loss can be interpreted as maximizing the lower bound of mutual information between the anchor and the positive example(oord2019representationlearningcontrastivepredictive; wu2020mutualinformationcontrastivelearning):
@@
$\log\left(1+\exp\left(\frac{x_{t+1}^{2}-x_{t+1}(x_{t}+\sigma)}{(x_{t}+\sigma)^{2}}h_{t}^{2}\right)\right)$
@@
$\log\left(1+\exp\left(\frac{h_{t}^{2}x_{t+1}^{4}}{(x_{t}+\sigma)^{3}}+\frac{h_{t}^{2}}{x_{t}+\sigma}\right)\right)$
@@
Appendix D Additional Experimental Results
D.1 Comprehensive Results of Comparison with Temporal Model
@@
D.2 Statistical Significance and Correlation Analysis of Focus Ratio and Memory Entropy
Pearson Correlation $r$ $p$-value
MSE & FR -0.5096 $3.1278\times 10^{-21}$
MSE & ME -0.8000 $4.4317\times 10^{-68}$
Table 13: Pearson Correlation Coefficients and $p$-values
We sampled Mean Squared Error (MSE), Focus Ratio (FR), and Memory Entropy (ME) at various epochs during multiple training iterations, resulting in a dataset comprising 300 sample points. Based on this data, we conducted statistical significance tests for FR and ME, as well as analyzed their correlations with MSE.
The analysis reveals significant negative correlations between Mean Squared Error (MSE) and both Focus Ratio (FR) and Memory Entropy (ME). Specifically, the Pearson correlation coefficient between MSE and FR is r=0.5096r=-0.5096 ($p<0.001$), indicating a moderate inverse relationship. Similarly, the correlation between MSE and ME is r=0.8000r=-0.8000 ($p<0.001$), demonstrating a robust negative association. These findings suggest that higher Focus Ratio and lower Memory Entropy are closely linked to reduced prediction errors, highlighting the critical role of attentional focus and efficient memory processes in enhancing predictive performance. The results underscore the importance of optimizing cognitive mechanisms, such as improving attention allocation and refining memory structures, to achieve greater accuracy in predictive tasks.
D.3 Detail Comparison of Improvements
To demonstrate that pre-training Mamba blocks with RCL can effectively enhance the temporal prediction capabilities of Mamba-based models, we present the performance improvements of four Mamba-based models after using pre-trained parameters. We conducted extensive tesWting on six datasets, each with an input length of 96 and prediction lengths of $\{96,192,336,720\}$. To clearly illustrate the performance improvements, we provide the percentage increase in MSE and MAE when using pre-trained parameters compared to not using them, as shown by the up-rate in Table 15.